-5x+y=-3
3x-8y=24

Answer by mananth(12270)

-5 x + 1 y = -3 .............1

3 x -8 y = 24 .............2
Eliminate y
multiply (1)by 8
Multiply (2) by 1
-40 x 8 y = -24
3 x -8 y = 24
Add the two equations
-37 x = 0
/ -37
x = 0
plug value of x in (1)
-5 x + 1 y = -3
0 + y = -3
y = -3